OpenStudy (anonymous):

If f(x) = the integral from -2 to x^4 of sqrt(t^2+8)dt then f'(x)=? I will write this as an equation as well below

OpenStudy (anonymous):

use main theorem of calculus again. Just substitute t^2 by (x^4)^2

OpenStudy (anonymous):

so answer should be sqr(x^8+8)

OpenStudy (anonymous):

If \[f(x)=\int\limits_{-2}^{x^4} \sqrt{t^2+8}dt \] f'(x)?

OpenStudy (anonymous):

For some reason it is saying that \[\sqrt{x^8+8}\] is the wrong answer

OpenStudy (anonymous):

sqr(x^8+8) isn't the right answer